OpenStudy (anonymous):

Two organ pipes, open at one end but closed at the other, are each 1.25 m long. One is now lengthened by 1.80 cm . Find the frequency of the beat they produce when playing together in their fundamental.

OpenStudy (anonymous):

Help guys! Please!

OpenStudy (anonymous):

Obfiusly, we have to find the fundamentals first. Remember the formula?

OpenStudy (anonymous):

 for fundamental resonance, L must be equal to 1/4 of lambda. From here, you can find lambda (and then frequency) of each pipe.

OpenStudy (anonymous):

So, L is 5m?

OpenStudy (anonymous):

No, no, you seem to missed the point. Here: 1/4 lambda =L lambda=4*L from this lambda, you can find frequency: lambda=v/f (v is speed of sound in air)

OpenStudy (anonymous):

lambda = 5m frequency = 5/344 m = .0145 is that good so far?

OpenStudy (anonymous):

actually, f=v/lambda :)

OpenStudy (anonymous):

whoops. so 344 m/s / 5m = 68.8 what do i do now?

OpenStudy (anonymous):

Yep, good. Now find the frequency of second pipe.

OpenStudy (anonymous):

344m/s / .072 m = 4777.78

OpenStudy (anonymous):

Correct. The result's quite surprising, because huge difference in length. Remember the definition of beat?

OpenStudy (anonymous):

unfortunately, i don't

OpenStudy (anonymous):

Okay, I'll be patiently teach you :) Beat is difference frequency. It occurs when two slightly different frequency played simultaneusly. Try pressing two neighboring piano button, you'll recognize it.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Can you find the frequency of the beat now?

OpenStudy (anonymous):

is the frequency just the difference of the two frequencies i found? so is it 4708.98?

OpenStudy (anonymous):

Yeah, mathematically, you're correct. But practically, beat will only occur (or realized) if the frequencies are slightly different, more like 400 Hz and 402 Hz.

OpenStudy (anonymous):

so, my answer is wrong then?

OpenStudy (anonymous):

No, you're right. But the problem must be reconsidered. Maybe the unit of second pipe shouldn't be in cm, but in m.

OpenStudy (anonymous):

I tried that and got -21.02 but the homework is still saying that i'm not getting the answer right???

OpenStudy (anonymous):

21,2 Hz seems more plausible. But you're in the right track if you explain in your homework that beat doesn't really occur.

OpenStudy (anonymous):

The homework is masteringphysics though, so I can't exactly say that through an online homework